# Adversary Resilient Learned Bloom Filters

### v1

@@ 5.3 Secure Classical Bloom Filter
$\alpha_{P}\text{FPR}(S^{\prime\prime},m_{A})+\alpha_{N}\text{FPR}(S^{\prime},m_{B})+(1-\alpha_{P}-\alpha_{N})\text{FPR}_{DB}(S,\mathcal{D},M)<\text{FPR}(S,m_{L}+m_{A}+m_{B}+\lambda)$
where $S^{\prime}\subset S$ is the set of elements in $S$ for which the Learned Oracle $L$ returns negative, $S^{\prime\prime}\subset S$ is the set of elements in $S$ for which the Learned Oracle $L$ returns positive, and where the probability is taken over the random coins of the pseudo-random permutations, the random coins used in the construction of the Learned Oracle, the random coins used in the construction of Backup Bloom Filters $A$ and $B$, the random coins used in the construction of the Secure Classical Bloom Filter, and the random coins used in the generation of the non-adversarial queries.
Proof.
The proof follows directly from the expression derived for the expected false positive probability of the Downtown Bodega Filter construction in Theorem 5.2 and the expression for the expected false positive probability of the Secure Classical Bloom Filter. ∎
@@ 5.4 Realistic Example
Mitzenmacher Broder [7] show that the false positive probability for a Classical Bloom Filter with $m$ bits encoding a set $S$, using $k(S,m)$ hash functions is
@@
In our analysis, we use the value of the number of hash functions $k(S,m)$ is always optimally chosen to be $k(S,m)=\ln 2\cdot(m/|S|)$ (this optimal value is also derived by Mitzenmacher Broder [7]). We further model the false positive rate of a Learned Oracle as being the same as the false positive rate of a Classical Bloom Filter encoding set $S$ but with a better false positive probability for the same memory budget. This is consistent with the assumptions made by prior work including Kraska et al [1] and Mitzenmacher [8].
@@
We note that the true negative probability of a Learned Oracle is merely the probability of a negative entry (which is constant as we are assuming set $S$ is constant) is not marked as a false positive. Let $Q_{N}$ be the fraction of true negative non-adversarial queries. We have
@@
To evaluate how much lower the false positive probability of a Learned Oracle needs to be for the Downtown Bodega Filter to perform better than the Secure Classical Bloom Filter, we may then use these derivations in Theorem 5.3.
We choose realistic values for our example from prior work on evaluating Learned Bloom Filters [1] on Google’s transparency report. We pick $2$ Megabytes as our memory budget, $m$, chosen from the range of values in Figure 10 of Kraska et al [1]. We choose the cardinality of the set we want to encode, $|S|$, as $1.7$ million based on the number of unique URLs in Google’s transparency report evaluated in Kraska et al [1]. With a memory budget of $2$ Megabytes, Kraska et al [1] demonstrate that a Learned Bloom Filter has $0.25$ of the False Positive Ratio of a Classical Bloom Filter, hence we use that as our value for $c$.
@@
M Memory Budget 2 Megabytes
$m_{L}$ Memory budget for Learned Oracle 1 Megabytes
$m_{A}$ Memory budget for Backup Bloom Filter $A$ $0.5$ Megabytes
$m_{B}$ Memory budget for Backup Bloom Filter $B$ $0.5$ Megabytes
@@
c FPR of Learned Bloom Filter / FPR of Classical Bloom Filter 0.25
@@
We use $128$ bits as the size of our security parameter, $\lambda$. For the case of the Downtown Bodega Filter, we let the Learned Oracle take $1$ Megabytes, while dividing the remaining $1$ Megabytes equally between Backup Classical Bloom Filters $A$ and $B$. Backup Classical Bloom Filters $A$ and $B$ encode $S^{\prime\prime}$ and $S^{\prime}$ respectively, which are both subsets of the set $S$ (refer to Definition 3.8). Since the false positive probability is adversely impacted by a larger set cardinality when the memory budget and the number of hash functions are kept constant, we take $|S^{\prime\prime}|=|S^{\prime}|=|S|$ as a lower bound on the performance of the Downtown Bodega Filter (i.e an upper bound on its false positive rate).
We take $\alpha$ to be a variable ranging from 00 to $0.2$ equally divided between $\alpha_{P}$ and $\alpha_{N}$. Our chosen values are summarized in Table 1. Figure 6 shows the results of our calculations. As can be seen, when the adversary has access to less than a certain cutoff fraction of the workload, the Downtown Bodega Filter outperforms the Secure Classical Bloom Filter for the same memory budget. The source code for our analysis can be found in an anonymously hosted Google Colab notebook [15].
Figure 6: The False Positive Rate of the Downtown Bodega Filter compared to the Secure Classical Bloom Filter in the hybrid adversarial setting where the adversary chooses between 00 and $0.2$ of the workload.
@@ 6 Open Problems
Problem 1: Secure Constructions for the Learned Bloom Filter in the unsteady setting. While we have provided secure constructions for the Learned Bloom Filter, our results only hold for the steady setting, where the query algorithm $\tilde{Q}_{s}$ may not modify the internal representation of the Learned Bloom Filter. Naor Yogev [3] provide secure constructions for the Classical Bloom Filter in an unsteady settings. We leave the formulation of secure constructions for the Learned Bloom Filter in the unsteady setting as an open problem.
Problem 2: Secure Learned Bloom Filters under a stronger adversarial model. The adversarial model we consider for the Learned Bloom Filter is stronger than Naor Yogev [3] in that it allows an adversary access to the internal state of the Bloom Filter, but it is weaker than Clayton et al [11] and Filic et al [12] as it does not allow the adversary to insert entries in the Learned Bloom Filter. We leave the formulation of secure constructions under an adversarial model that allows insertion as an open problem.
Problem 3: Secure Learned Bloom Filters under computationally unbounded adversaries. In this work, we only consider probabilistic polynomial time (PPT) adversaries. An open problem is to prove or disprove any security guarantees for our constructions against computationally unbounded adversaries. Alternatively, an open problem is to provide secure constructions against computationally unbounded adversaries. Naor Yogev [3] discuss computationally unbounded adversaries in Section 5 of their work.

### v3

@@ 5.3 Secure Classical Bloom Filter
$\begin{split}\alpha_{P}\text{FPR}(S^{\prime\prime},m_{A})+\alpha_{N}\text{FPR}(S^{\prime},m_{B})+(1-\alpha_{P}-\alpha_{N})\text{FPR}_{DB}(S,\mathcal{D},M)\\ <\text{FPR}(S,m_{L}+m_{A}+m_{B}+\lambda)\end{split}$ (2)
where $S^{\prime}\subset S$ is the set of elements in $S$ for which the Bloom Filter Learning Model $L$ returns negative, $S^{\prime\prime}\subset S$ is the set of elements in $S$ for which the Bloom Filter Learning Model $L$ returns positive, and where the probability is taken over the random coins of the pseudo-random permutations, the random coins used in the construction of the Bloom Filter Learning Model, the random coins used in the construction of Backup Bloom Filters $A$ and $B$, the random coins used in the construction of the Secure Classical Bloom Filter, and the random coins used in the generation of the non-adversarial queries.
Proof 5.6.
The proof follows directly from the expression derived for the expected false positive probability of the Downtown Bodega Filter construction in Theorem 5.3 and the expression for the expected false positive probability of the Secure Classical Bloom Filter.
@@ 5.4 Realistic Example
Mitzenmacher Broder [MB04] show that the false positive probability for a Classical Bloom Filter with $m$ bits encoding a set $S$, using $k(S,m)$ hash functions is
@@
In our analysis, we use the value of the number of hash functions $k(S,m)$ is always optimally chosen to be $k(S,m)=\ln 2\cdot(m/|S|)$ (this optimal value is also derived by Mitzenmacher Broder [MB04]). We further model the false positive rate of a Bloom Filter Learning Model as being the same as the false positive rate of a Classical Bloom Filter encoding set $S$ but with a better false positive probability for the same memory budget. This is consistent with the assumptions made by prior work including Kraska et al [KBC+18] and Mitzenmacher [Mit18].
@@
We note that the true negative probability of a Bloom Filter Learning Model is merely the probability of a negative entry (which is constant as we are assuming set $S$ is constant) is not marked as a false positive. Let $Q_{N}$ be the fraction of true negative non-adversarial queries. We have
@@
To evaluate how much lower the false positive probability of a Bloom Filter Learning Model needs to be for the Downtown Bodega Filter to perform better than the Secure Classical Bloom Filter, we may then use these derivations in Theorem 5.5.
We choose realistic values for our example from prior work on evaluating Learned Bloom Filters [KBC+18] on Google’s transparency report. We pick $2$ Megabytes as our memory budget, $m$, chosen from the range of values in Figure 10 of Kraska et al [KBC+18]. We choose the cardinality of the set we want to encode, $|S|$, as $1.7$ million based on the number of unique URLs in Google’s transparency report evaluated in Kraska et al [KBC+18]. With a memory budget of $2$ Megabytes, Kraska et al [KBC+18] demonstrate that a Learned Bloom Filter has $0.25$ of the False Positive Ratio of a Classical Bloom Filter, hence we use that as our value for $c$. We assume $Q_{N}$ to be $0.5$ for this example (results for the complete range of values of $Q_{N}$ can be found in Appendix C).
@@
M Total memory budget 2 MB
$m_{L}$ Memory budget for Bloom Filter Learning Model 1 MB
$m_{A}$ Memory budget for Backup Classical Bloom Filter $A$ $0.5$ MB
$m_{B}$ Memory budget for Backup Classical Bloom Filter $B$ $0.5$ MB
@@
c $\frac{\text{FPR of Bloom Filter Learning Model})}{\text{FPR of Classical BloomFilter}}$ for same memory budget 0.25
@@
We use $128$ bits as the size of our security parameter, $\lambda$. For the case of the Downtown Bodega Filter, we let the Bloom Filter Learning Model take $1$ Megabytes, while dividing the remaining $1$ Megabytes equally between Backup Classical Bloom Filters $A$ and $B$. Backup Classical Bloom Filters $A$ and $B$ encode $S^{\prime\prime}$ and $S^{\prime}$ respectively, which are both subsets of the set $S$ (refer to Definition 3.4).
We take $\alpha$ to be a variable ranging from 00 to $1$ equally divided between $\alpha_{P}$ and $\alpha_{N}$ (refer to Appendix C for other strategies of partitioning $\alpha$ between $\alpha_{P}$ and $\alpha_{N}$). Our chosen values are summarized in Table 1. Figure 6 shows the results of our calculations. As can be seen, when the adversary has access to less than a certain cutoff fraction of the workload, the Downtown Bodega Filter outperforms the Secure Classical Bloom Filter for the same memory budget. The $C$ source code for our model and analysis can be found in an anonymously hosted code repository [Bod24]. A thorough experimental evaluation of the Hybrid Model across multiple datasets and a comprehensive range of model parameters can be found in Appendix C.
Figure 6: The FPR of the Downtown Bodega Filter compared to the Secure Classical Bloom Filter in the hybrid adversarial setting as evaluated on the Google transparency report.
@@ 6 Open Problems
Problem 1: Secure Constructions for the Learned Bloom Filter in the unsteady setting. While we have provided secure constructions for the Learned Bloom Filter, our results only hold for the steady setting, where the query algorithm $\tilde{Q}_{s}$ may not modify the internal representation of the Learned Bloom Filter. Naor Yogev [NE19] provide secure constructions for the Classical Bloom Filter in an unsteady settings. We leave the formulation of secure constructions for the Learned Bloom Filter in the unsteady setting as an open problem.
Problem 2: Secure Learned Bloom Filters under a stronger adversarial model. The adversarial model we consider for the Learned Bloom Filter is stronger than Naor Yogev [NE19] in that it allows an adversary access to the internal state of the Bloom Filter, but it is weaker than Clayton et al [CPS19] and Filic et al [FPUV22] as it does not allow the adversary to insert entries in the Learned Bloom Filter. We leave the formulation of secure constructions under an adversarial model that allows insertion as an open problem.
Problem 3: Secure Learned Bloom Filters under computationally unbounded adversaries. In this work, we only consider probabilistic polynomial time (PPT) adversaries. An open problem is to prove or disprove any security guarantees for our constructions against computationally unbounded adversaries. Alternatively, an open problem is to provide secure constructions against computationally unbounded adversaries. Naor Yogev [NE19] discuss computationally unbounded adversaries in Section 5 of their work.
@@ 7 Acknowledgements
Appendices
Appendix A contains discussion of an alternate secure construction of the Learned Bloom Filter which we call the Uptown Bodega Filter. Appendix B contains a brief but self-contained treatment of pseudo-random permutations. Appendix C contains detailed results across a wide range of model parameters for the Hybrid Model. Appendix D contains an overview of our open-source implementation of the Downtown Bodega Filter as well as the Secure Classical Bloom Filter. The appendix also contains results from experiments to investigate how the False Positive Rate of the Downtown Bodega Filter varies compared to the Secure Classical Bloom Filter as we vary the memory budget.
Appendix A The Uptown Bodega Filter
In addition to the Downtown Bodega Filter, discussed in the main text, we consider a second secure construction for the Learned Bloom Filter built on top of the Sandwiched Learned Bloom Filter and a pseudo-random permutation.
Definition A.1 (Uptown).
Theorem A.2.
Our proof follows the same line of reasoning as Theorem $4.8$ of Naor Yogev [NE19] which proves that, if pseudo-random permutations exist, any probabilistic polynomial time adversary cannot distinguish between the Classical Bloom Filter with pseudo-random permutations and Challenge 1 by more than a negligible advantage. Due to Lemma 3.2, this result also holds for the Uptown Bodega Filter. The crux of the proof relies on the fact that running $F_{k}$ on an adversary’s queries permits us to consider the queries as random and not chosen adaptively by an adversary, while having no effect on the correctness of the Uptown Bodega Filter.
Proof A.3.
We construct a Uptown Bodega Filter $UB_{k,r}$ from a Sandwiched Bloom Filter $SB_{r}$ as follows. Choose the first key $k\in{\{0,1\}}^{\lambda}$ from pseudo-random permutation, $F_{k}$, over ${\{0,1\}}^{\log|U|}$. Let $S_{k}=F_{k}(S)={\{F_{k}(x):x\in S\}}$. We initialize $SB_{R}$ with $S_{k}$. For each input $x$, our query algorithm outputs $SB_{k,r}(F_{k}(x))$. Just like in the classical Bloom Filter construction of [NE19], the only additional memory required is to store the key $k$ for $F_{k}$, which is $\lambda$ bits. Similarly, the running time of the query algorithm of the Uptown Bodega Filter is one pseudo-random permutation more than the running time of the query algorithm of the Sandwiched Learned Bloom Filter.
Now we show that no probabilistic polynomial time (PPT) adversary $A$ can distinguish between the Uptown Bodega Filter we constructed using $R(\cdot)$ and the Uptown Bodega Filter construction that uses the pseudo-random permutation $F_{k}$ by more than a negligible advantage. Suppose that there does exist a non-negligible function $\delta(\lambda)$ such that $A$ can attack $UB_{k,r}$ and find a false positive with probability $\epsilon+\delta(\lambda)$. We can run $A$ on $UB_{k,r}$ where the oracle is replaced by an oracle that is either random or pseudo-random. We return $1$ if $A$ successfully finds a false positive. This implies that we may distinguish between a truly random permutation and a pseudo-random permutation with probability $\geq\delta(\lambda)$. This contradicts the indistinguishability of pseudo-random permutations.
A.1 Utility of the Uptown Bodega Filter
While we have shown the existence of $(n,\epsilon)$-adversarial resilient Uptown Bodega Filters, it is unclear whether there is any utility in using an Uptown Bodega Filter over a Downtown Bodega Filter or a Secure Classical Bloom Filter. This is due to the fact that we train our learning model on $F_{k}(x)\in S$, a pseudo-random permutation of the set $S$, instead of training it directly on $S$. Intuitively, any structure in $S$ that can be “learned” by a Bloom Filter Learning Model is destroyed by taking the pseudo-random permutation of $S$. In other words, if pseudo-random permutations exist, then to any Bloom Filter Learning Model that runs in Probabilistic Polynomial Time (PPT), $F_{k}(x)\in S$ should appear the same as $|S|$ values sampled from a truly random oracle $R(\cdot)$. A precise proof for this is beyond the scope of this paper.
Appendix B Pseudo-random Permutations
This section provides a brief self-contained treatment of pseudo-random permutations adapted from Chapter 3 of Katz Lindell [KL14] here.
Let $\text{Perm}_{n}$ be the set of all permutations on $\{0,1\}^{n}$.
Definition B.1.
Definition B.2.
Definition B.3.
Appendix C Hybrid Model - Additional Experiments
This section contains additional experiments under the Hybrid Adversarial Model setting described in Section 5. We analyze the Hybrid Model using $494$ lines of C code in the model/ directory of our open-source implementation [Bod24].
C.1 Varying Non-Adversarial True Negatives ($Q_{N}$)
In the main text, we assumed the fraction of non-adversarial queries that were true negatives, $Q_{N}$ to be $0.5$. Here we show the results for the entire range of values of $Q_{N}\in[0,1]$ for $\alpha$ taking $4$ values: $0.2,0.3,0.5,1.0$, with each value partitioned equally between $\alpha_{P}$ and $\alpha_{N}$. The results are in Figure 7.
(a) $\alpha=0.2$
(b) $\alpha=0.3$
(c) $\alpha=0.5$
(d) $\alpha=1.0$
Figure 7: The FPR of the Downtown Bodega Filter compared to the Secure Classical Bloom Filter in the hybrid adversarial setting as evaluated on the Google transparency report with $Q_{N}$ taking a range of values in the interval $[0,1]$
Figure 8: The FPR of the Downtown Bodega Filter compared to the Secure Classical Bloom Filter in the hybrid adversarial setting as evaluated on the Google transparency report as we vary the fraction of adversarial queries that generate false positives
C.2 Adversary Strategies
In the main text, we explored the case where the adversary divides $\alpha N$ queries equally between queries that generate False Positives, and queries that generate False Negatives. We conduct an experiment here for all partitions of $\alpha$ between $\alpha_{P}$ and $\alpha_{N}$ for $\alpha=0.2$ to see how the False Positive Rate of the Downtown Bodega Filter is impacted. We vary the fraction of $\alpha$ assigned to $\alpha_{P}$ from 00 to $1$. Here, 00 means the adversary spends their entire budget of $\alpha N$ queries on true negatives, and $1$ means the adversary spends their entire budget of $\alpha N$ queries on false positives. Our earlier experiments are setting this fraction to $0.5$, in this framework. The results are in Figure 8
C.3 Additional Datasets
We conduct experiments on two other common evaluation datasets used in prior work on the Learned Bloom Filter [SM24, VKMK21, DS20]. We include a brief description of these datasets, adapted from Sato and Matsui’s recent NeurIPS 2023 work on Fast Partitioned Learned Bloom Filters [SM24].
  • Malicious URLs Dataset [mal]: The URLs dataset contains $223,088$ malicious and $428,118$ benign URLs.
  • EMBER Dataset [AR18]: This dataset contains $300,000$ malicious and $400,000$ benign files.
For these datasets, we change the values of the cardinality of the set to encode, $n$, in Table 1. For all other model parameters, we use the same ones listed in the table. Figures 9 and  10 show the results for the Malicious URLs Dataset and the EMBER dataset respectively.
Figure 9: The FPR of the Downtown Bodega Filter compared to the Secure Classical Bloom Filter in the hybrid adversarial setting as evaluated on the Malicious URLs dataset.
Figure 10: The FPR of the Downtown Bodega Filter compared to the Secure Classical Bloom Filter in the hybrid adversarial setting as evaluated on the EMBER dataset.
Appendix D Implementation
We implement both the Downtown Bodega Filter and the Secure Classical Bloom Filter in $441$ lines of Python 3 code, which can be found in the src/ directory of our open-source implementation [Bod24]. Similar to Naor Yogev [NE19], we implement pseudo-random permutations using AES with a block size of $128$ bits. We use CBC-mode. Our implementation makes use of the PyCryptoDome [PyC] library for these cryptographic primitives. Our implementation is modular and allows any classification model to be easily plugged in. We tested the correctness of our implementation on a broad range of common classification models including the Random Forest model, Gaussian Naive Bayes, the Gradient Boosting Classifier, Support Vector Machine based Classifiers, and Adaptive Boosting. We relied on implementations of these models provided by scikit-learn.
D.1 Experimental Setup
Similar to prior work [SM24, VKMK21], we tested our implementation on the Malicious URLs dataset [mal], which we describe in Appendix C. We unoformly randomly sample $20\%$ of the dataset to reduce our experimentation time. We use a set of $20$ features to train our Learning Models including: length of the URL, whether the URL contains an IP address, whether the URL uses a shortening service, whether the URL is “abnormal”, the digit count and letter count of the URL, and whether the URL contains special symbols such as @,?, * etc. This set of features for the Malicious URLs Dataset is common in open-source learning models, and similar features have been used by Vaidiya, Knorr, Mitzenmacher et al in prior work [VKMK21].
D.2 Methodology
We calculate the False Positive Rate of both the Secure Classical Bloom Filter as well as the Downtown Bodega Filter by uniformly randomly sampling URLs that are not malicious from Malicious URLs Dataset and counting how many of them are returned as false positive by both filters. We calculate the memory used by the learning model by using joblib to create a persistent version [Sci] of the model, and measuring the size of the persistent version of the model. After subtracting the memory used by the learning model from the memory budget, we divide the remaining memory budget equally between Backup Classical Bloom Filter $A$ and Backup Classical Bloom Filter $B$.
D.3 Results
Figure 11 shows how how the False Positive Rate varies for both the Secure Classical Bloom Filter as well as the Downtown Bodega Filter as we modify the memory budget available. This figure uses the Gaussian Naive-Bayes Classifier as the Bloom Filter Learning Model. Figure 12 shows the same results using a Linear Support Vector Classifier as the BLoom Filter Learning Model. [^1] All the data used to plot these results can be found in src/results_plot.py in our code repository.
Figure 11: The FPR of the Downtown Bodega Filter using a Gaussian Naive-Bayes Classifier as its Bloom Filter Learning Model compared to the Secure Classical Bloom Filter with varying memory budget as evaluated on the Malicious URLs Dataset
Figure 12: The FPR of the Downtown Bodega Filter using a Linear Support Vector Classifier as its Bloom Filter Learning Model compared to the Secure Classical Bloom Filter with varying memory budget as evaluated on the Malicious URLs Dataset